# Dynamic-SUPERB Phase-2: A Collaboratively Expanding Benchmark for Measuring the Capabilities of Spoken Language Models with 180 Tasks

### v1

@@ Appendix B Dynamic-SUPERB tasks
| Speech / Speaker & Language / Speaker / Speaker Recognition | SuperbSD | Libri2Mix Test (Cosentino et al., 2020) |
| Speech / Speaker & Language / Speaker / Speaker Recognition | SuperbSV | VoxCeleb1 Test (Nagrani et al., 2020) |
@@
| Speech / Spoken Language Understanding / Speech Translation | SuperbST | CoVoST2 Test (Wang et al., 2021) |
| Speech / Phonetics, Phonology, Prosody / Prosody/Prosodic Feature Classification | Stress Detection | MIRSD (CHEN & JANG, 2012) |
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Phoneme Recognition Tasks | Superb PR | LibriSpeech-TestClean (Panayotov et al., 2015) |
| Speech / Safety and Security / Spoofing and Anti-Spoofing | Spoof Detection | ASVspoof2015 (Wu et al., 2015) |
| Speech / Safety and Security / Spoofing and Anti-Spoofing | Spoof Detection | ASVspoof2017 (Delgado et al., 2018) |
| Speech / Spoken Language Understanding / Linguistic Structure & Grammar | PoS Estimation | LibriTTS_PoS (Zen et al., 2019) |
| Speech / Spoken Language Understanding / Linguistic Structure & Grammar | PoS Estimation | LibriTTS_PoS_with_transcription (Zen et al., 2019) |
| Speech / Spoken Language Understanding / Linguistic Structure & Grammar | Code Switching Semantic Grammar Acceptability Comparison | CSZS-zh-en (Huang et al., 2024b) |
@@
| Speech / Speech Recognition / Language | Code-switching Speech Recognition | NTUML2021 (Yang et al., 2024a) |
| Speech / Speech Recognition / Language | PTBR Speech Recognition | CORAA (Candido Junior et al., 2022) |
@@
| Speech / Speech Recognition / Language | Superb OODAsr Ar | CommonVoice7-Test (Ardila et al., 2020) |
| Speech / Speech Recognition / Language | Superb OODAsr Es | CommonVoice7-Test (Ardila et al., 2020) |
| Speech / Speech Recognition / Language | Superb OODAsr Zh | CommonVoice7-Test (Ardila et al., 2020) |
@@
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | LJSpeech (Ito & Johnson, 2017), RirsNoises-SmallRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | LJSpeech (Ito & Johnson, 2017), RirsNoises-MediumRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | LJSpeech (Ito & Johnson, 2017), RirsNoises-LargeRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | VCTK (Yamagishi et al., 2019), RirsNoises-SmallRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | VCTK (Yamagishi et al., 2019), RirsNoises-MediumRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | VCTK (Yamagishi et al., 2019), RirsNoises-LargeRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019), MUSAN-Music (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019), MUSAN-Speech (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019), MUSAN-Gaussian (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019), MUSAN-Noise (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019), MUSAN-Music (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019), MUSAN-Speech (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019), MUSAN-Gaussian (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019), MUSAN-Noise (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017), MUSAN-Music (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017), MUSAN-Speech (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017), MUSAN-Gaussian (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017), MUSAN-Noise (Snyder et al., 2015) |
@@
| Speech / Spoken Language Understanding / Intent & Meaning | Sentiment Analysis | SLUE-VoxCeleb (Shon et al., 2022) |
@@
| Speech / Spoken Language Understanding / Intent & Meaning | Superb IC | FluentSpeechCommands-Test (Lugosch et al., 2019) |
| Speech / Spoken Language Understanding / Intent & Meaning | Superb SF | AudioSnips-Test (Yang et al., 2021) |
| Speech / Spoken Language Understanding / Intent & Meaning | Intent Classification | FluentSpeechCommands-Action (Lugosch et al., 2019) |
| Speech / Spoken Language Understanding / Intent & Meaning | Intent Classification | FluentSpeechCommands-Location (Lugosch et al., 2019) |
| Speech / Spoken Language Understanding / Intent & Meaning | Intent Classification | FluentSpeechCommands-Object (Lugosch et al., 2019) |
| Speech / Spoken Language Understanding / Intent & Meaning | Sentence Grammar Acceptability | sBLIMP (Nguyen et al., 2020) |
| Speech / Paralinguistics / Emotion Analysis | Superb ER | Iemocap (Busso et al., 2008) |
| Speech / Paralinguistics / Emotion Analysis | Emotion Recognition | Multimodal EmotionLines Dataset (Poria et al., 2019) |
| Speech / Paralinguistics / Vocal Event Detection | Human Non-Speech Sound Recognition | Nonspeech7k-test (Rashid et al., 2023), CommonVoice-DeltaSegment-15 (Ardila et al., 2020) |
| Speech / Phonetics, Phonology, Prosody / Prosody / Prosodic Quality Assessment | L2 English Accuracy | speechocean762-Ranking (Zhang et al., 2021) |
| Speech / Phonetics, Phonology, Prosody / Prosody / Prosodic Quality Assessment | L2 English Accuracy | speechocean762-scoring (Zhang et al., 2021) |
| Speech / Phonetics, Phonology, Prosody / Prosody / Prosodic Quality Assessment | Heteronym Differentiation | HeteronymEn (Synthesized heteronym pairs from Wikipedia HeteroEnWiki with the OpenAI text-to-speech service OpenAI (2024)) |
| Speech / Phonetics, Phonology, Prosody / Prosody / Prosodic Quality Assessment | Multilingual Pronunciation Similarity | VoxAngeles (Chodroff et al., 2024) |
| Speech / Safety and Security / Synthetic Speech Detection | DeepFake Voice Recognition | DEEP-VOICE (Bird & Lotfi, 2023) |
| Speech / Safety and Security / Synthetic Speech Detection | Fraud Robocall Recognition | CallHome (Katerenchuk et al., 2018) |
| Speech / Safety and Security / Synthetic Speech Detection | Fraud Robocall Recognition | Promo (Generated using XTTS-v1 Casanova et al. (2024) on the Robocall Audio Dataset Prasad & Reaves ) |
| Speech / Safety and Security / Synthetic Speech Detection | Fraud Robocall Recognition | Robocall (Prasad & Reaves, ) |
| Speech / Safety and Security / Synthetic Speech Detection | Enhancement Detection | LibriTTS-TestClean (Zen et al., 2019), WHAM (Wichern et al., 2019) |
| Speech / Speech Recognition / Specific Recognition Tasks | Superb KS | SpeechCommandsV1 Test (Warden, 2018) |
| Speech / Speech Recognition / Specific Recognition Tasks | Superb QbE | Quesst14 Eval (Anguera et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | HEAR Spoken Commands Classification | SpeechCommands-5hr (Warden, 2018) |
| Speech / Speech Recognition / Specific Recognition Tasks | Spoken Term Detection | LibriSpeech TestOther (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Spoken Term Detection | LibriSpeech TestClean (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Multi Speaker Detection | LibriSpeech TestClean (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Speech Text Matching | LibriSpeech TestClean (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Speech Text Matching | LibriSpeech TestOther (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Speech Command Recognition | GoogleSpeechCommandsV1 (Warden, 2018) |
| Speech / Speech Recognition / Spontaneous Speech | Superb OOD Asr Spon | SBCSAE Test (Du Bois et al., 2000 – 2005) |
| Speech / Speaker & Language / Speaker / Speaker Characteristics | Age Classification | CommonVoiceCorpus Test (Ardila et al., 2020) |

### v2

@@ Appendix B Dynamic-SUPERB tasks
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Pronounciation Evaluation | Heteronym Differentiation | HeteronymEn (self-constructed) |
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Pronounciation Evaluation | L2 English Accuracy | speechocean762-Ranking (Zhang et al., 2021) |
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Pronounciation Evaluation | L2 English Accuracy | speechocean762-scoring (Zhang et al., 2021) |
| Speech / Phonetics, Phonology, Prosody / Phonetics and Phoneme Processing / Pronounciation Evaluation | Multilingual Pronunciation Similarity | VoxAngeles (Chodroff et al., 2024) |
| Speech / Phonetics, Phonology, Prosody / Prosody / Prosodic Feature Classification | Stress Detection | MIRSD (CHEN & JANG, 2012; Tseng, 2008; Tsai, 2015) |
@@
| Speech / Safety & Security / Spoofing and Anti-Spoofing | Spoof Detection | ASVspoof2015 (Wu et al., 2015) |
| Speech / Safety & Security / Spoofing and Anti-Spoofing | Spoof Detection | ASVspoof2017 (Delgado et al., 2018) |
| Speech / Safety & Security / Synthetic Speech Detection | Deep Fake Voice Recognition | DEEP-VOICE (Bird & Lotfi, 2023) |
| Speech / Safety & Security / Synthetic Speech Detection | Enhancement Detection | LibriTTS-TestClean (Zen et al., 2019) WHAM (Wichern et al., 2019) |
| Speech / Safety & Security / Synthetic Speech Detection | Fraud Robocall Recognition | CallHome (Katerenchuk et al., 2018) |
| Speech / Safety & Security / Synthetic Speech Detection | Fraud Robocall Recognition | Promo (self-constructed) |
| Speech / Safety & Security / Synthetic Speech Detection | Fraud Robocall Recognition | Robocall Prasad & Reaves |
| Speech / Speaker & Language / Speaker / Speaker Characteristics | Age Classification | CommonVoiceCorpus-Test (Ardila et al., 2020) |
@@
| Speech / Speaker & Language / Speaker / Speaker Recognition | Speaker Verification | LibriSpeech-TestOther (Panayotov et al., 2015) |
@@
| Speech / Speaker & Language / Speaker / Speaker Recognition | SUPERB Speaker Diarization | Libri2Mix-Test (Cosentino et al., 2020) |
| Speech / Speaker & Language / Speaker / Speaker Recognition | SUPERB Speaker Verification | SUPERB Hidden Set (self-reconstructed) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017) MUSAN-Gaussian (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017) MUSAN-Music (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017) MUSAN-Noise (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | LJSpeech (Ito & Johnson, 2017) MUSAN-Speech (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019) MUSAN-Music (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019) MUSAN-Noise (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019) MUSAN-Speech (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise Detection | VCTK (Yamagishi et al., 2019) MUSAN-Gaussian (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019) MUSAN-Gaussian (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019) MUSAN-Music (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019) MUSAN-Noise (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Noise SNR Level Prediction | VCTK (Yamagishi et al., 2019) MUSAN-Speech (Snyder et al., 2015) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | LJSpeech (Ito & Johnson, 2017) RirsNoises-LargeRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | LJSpeech (Ito & Johnson, 2017) RirsNoises-MediumRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | LJSpeech (Ito & Johnson, 2017) RirsNoises-SmallRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | VCTK (Yamagishi et al., 2019) RirsNoises-LargeRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | VCTK (Yamagishi et al., 2019) RirsNoises-MediumRoom (Ko et al., 2017) |
| Speech / Speech Enhancement / Degradation Detection | Reverberation Detection | VCTK (Yamagishi et al., 2019) RirsNoises-SmallRoom (Ko et al., 2017) |
| Speech / Speech Recognition / Language | Code-switch Speech Recognition | NTUML2021 (Yang et al., 2024a) |
@@
| Speech / Speech Recognition / Language | PTBR Speech Recognition | CommonVoice17-Test (Ardila et al., 2020) |
@@
| Speech / Speech Recognition / Language | ASR | LibriSpeech-TestOther (Panayotov et al., 2015) |
| Speech / Speech Recognition / Language | SUPERB OOD ASR (Ar) | CommonVoice7-Test (Ardila et al., 2020) |
| Speech / Speech Recognition / Language | SUPERB OOD ASR (Es) | CommonVoice7-Test (Ardila et al., 2020) |
| Speech / Speech Recognition / Language | SUPERB OOD ASR (Spontaneous) | CHIME6-Test (Barker et al., 2018; Watanabe et al., 2020) |
| Speech / Speech Recognition / Language | SUPERB OOD ASR (Zh) | CommonVoice7-Test (Ardila et al., 2020) |
@@
| Speech / Speech Recognition / Specific Recognition Tasks | Multi-Speaker Detection | LibriSpeech-TestClean (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Speech Text Matching | LibriSpeech-TestClean (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Speech Text Matching | LibriSpeech-TestOther (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Spoken Term Detection | LibriSpeech-TestClean (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | Spoken Term Detection | LibriSpeech-TestOther (Panayotov et al., 2015) |
| Speech / Speech Recognition / Specific Recognition Tasks | SUPERB Keyword Spotting | SpeechCommandsV1-Test (Warden, 2018) |
| Speech / Speech Recognition / Specific Recognition Tasks | SUPERB Query by Example | Quesst14-Eval (Anguera et al., 2015) |
| Speech / Spoken Language Understanding / Intent & Meaning | SUPERB Intent Classification | SLURP (Bastianelli et al., 2020) |
| Speech / Spoken Language Understanding / Intent & Meaning | SUPERB Intent Classification (Action) | SLURP-Action (Bastianelli et al., 2020) |
| Speech / Spoken Language Understanding / Intent & Meaning | SUPERB Intent Classification (Intent) | SLURP-Intent (Bastianelli et al., 2020) |
@@
| Speech / Spoken Language Understanding / Intent & Meaning | Speech Sentiment Analysis | MELD (Poria et al., 2019) |
@@
| Speech / Spoken Language Understanding / Intent & Meaning | SUPERB Slot Filling | AudioSnips-Test (Yang et al., 2021) |
| Speech / Spoken Language Understanding / Linguistic Structure & Grammar | Code-Switching Semantic Grammar Acceptability Comparison | CSZS-zh-en (Huang et al., 2024b) |
| Speech / Spoken Language Understanding / Linguistic Structure & Grammar | PoS Estimation LibriTTS PoS | LibriTTS (Zen et al., 2019) |
| Speech / Spoken Language Understanding / Linguistic Structure & Grammar | PoS Estimation LibriTTS PoS with transcription | LibriTTS (Zen et al., 2019) |
| Speech / Spoken Language Understanding / Speech Translation | SUPERB Speech Translation | CoVoST2-Test (Wang et al., 2021) |